# Predicting Field Experiments with Large Language Models

### v1

@@ 4.3 Prediction
In the final stage, our framework takes a field experiment’s conclusion and its two variants, and experiment settings as input and generates two parallel prediction prompts: Basic Prediction Prompt and Chain-of-Thought (CoT) Prediction Prompt, which are then prompted to GPT to get predictions by asking GPT to select one conclusion from three conclusions. Since CoT has proven to be capable of improving general performance of LLMs on downstream tasks Wei et al. (2022), it’s interesting to see how its effect on this prediction scenario.
Figure 6: Basic Prediction Prompt
Figure 7: Chain-of-Thought (CoT) Prediction Prompt
@@
The basic prediction prompt is shown in Figure 6. It consists of a background information section, a question section, and necessary instructions. Specifically, the background information section contains the general goal of the experiment (such as exploring the impact of job training on income), treatments (such as receiving job training or not), experiment duration (such as seven weeks), outcomes (such as income), participant information (such as people seeking jobs in New England), and experiment workflow (such as when and how training was given and outcomes were recorded). All of these were extracted from a target paper, which is the same as bullet points “A” to “F” from Figure 4.
@@
Following the background information, a question section is automatically generated and entered into the templates. Specifically, the original conclusion, its reversed variant, and its unrelated variant are shuffled and substitute the placeholders “option 1”, “option 2”, and “option 3”. Meanwhile, instructions tell GPT that the prediction of conclusions is under the context of the field experimental settings, asking GPT to choose one option from the three options as it deems correct.
@@
As we are also interested in how CoT would improve such prediction, Figure 7 shows the CoT Prediction Prompt, which follows a similar logic as the basic prediction but integrates CoT strategies to boost the performance Wei et al. (2022). Initially, the framework prompts the experiment settings and three options of a conclusion to GPT, instructing it to think about decisive elements that help choose the correct option. Upon receiving the decisive elements from GPT, the framework prompts GPT to make the selection among three options to get a predicted conclusion. It is worth noting that the entire process is within the same GPT session for a conclusion.
@@
Although either prompt strategy generates a prediction for a given input, LLMs are stochastic models, meaning that their responses may vary to the same prompt. To handle such randomness in experiment simulations, Leng & Yuan Leng and Yuan (2023) set the temperature to 0 and always get fixed responses from synthetic participants of lab experiments, which is a strategy to eliminate the stochasticity of LLMs totally. By contrast, Brand et al. Brand et al. (2023) repeated the same prompt 300 times and use the averaged number as the result of Willingness-to-Pay from customers role-played by LLMs. Here, we take the latter approach by incorporating the stochasticity of LLM outputs since this stochasticity of LLMs is similar to how the same human participant might respond differently when presented with the same instruction Demszky et al. (2023). Specifically, we repeat the same prediction prompt several times and calculate an average accuracy as the final result. For example, if the framework is running the basic prediction, a filled-out prompt based on Figure 6 will be repeated a given number of times to get a stable result. The determination of a proper repeat number will be further discussed in the Results section.
@@
Parameter-wise, no fine-tuning is involved in any stage of the proposed framework, and all parameters of OpenAI API and Anthropic API are set to default. Whereas Horton Horton (2023) harnessed fine-tuned LLMs in lab experimental simulation to better follow instructions, Coda-Forno et al. Coda-Forno et al. (2024) simulated several human behaviors by LLMs without fine-tuning. The use of fine-tuned LLMs complicates the reproduction attempts since other researchers don’t have access to the same model in the existing papers Bail (2024). Additionally, avoiding fine-tuning LLMs saves computing resources and mitigates environmental impacts Ray (2023), especially when the pretraining of LLMs is enough to make them capable of downstream tasks Liu et al. (2023).
5 Results
@@
In this section, we test the framework on 319 field experiments that contain a total of 1638 conclusions and discuss the performance.
@@
“Conclusion Accuracy” and “Paper Accuracy” are used to evaluate the prediction performance under different settings. As shown in Figure 3, the framework generates a prompt for each conclusion. The generated prompt either follows the template in Figure 6 or 7, depending on which prompt strategy (basic or CoT) the framework applies. Each prompt instructs GPT to output a predicted conclusion. If the predicted conclusion matches the true conclusion, that attempt is counted as correct. As aforementioned, the framework repeats such an attempt for a set number of times for each conclusion to get a stable result. Therefore, we define “Conclusion Accuracy” as the percentage of correct predictions among a set number of attempts.
@@
Conclusion Accuracy=Number of correct predictionsTotal number of $\text{Conclusion Accuracy}=\frac{\text{Number of correct predictions}}{\text{Total number of predictions}}\times 100\%$
Given that a field experiment may contain multiple conclusions, “Paper Accuracy” is the average of all “Conclusion Accuracy” within a paper.
Paper $\text{Paper Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\text{Conclusion Accuracy}_{i}$
5.1 Prediction Performance Overview
GPT4-turbo Basic CoT
Conclusion Accuracy Paper Accuracy Conclusion Accuracy Paper Accuracy
Repeat = 10 62% 63% 73% 73%
Repeat = 20 61% 62% 73% 74%
Repeat = 30 61% 62% 73% 74%
Sample Size: 1638 conclusions from 319 papers
Table 1: Prediction Accuracy under Different Repeats
Repeat = 20 Basic CoT
Conclusion Accuracy Paper Accuracy Conclusion Accuracy Paper Accuracy
GPT3-turbo 60% 61% 66% 67%
GPT4-turbo 61% 62% 73% 74%
GPT4o 73% 74% 76% 77%
Sample Size: 1638 conclusions from 319 papers
Table 2: Prediction Accuracy under 20 Repeats by GPT Models
@@ 5.1 Prediction Performance Overview
Table 1 reports two types of accuracy for both strategies under different repeats by GPT4-turbo (gpt-4-turbo-2024-04-09). The best results for the basic strategy are obtained under 10 repeats, which are 62% for conclusion accuracy and 63% for paper accuracy. By comparison, the best CoT results are 73% and 74% under either 20 or 30 repeats, which are generally 17% higher than basic results and aligns with prior literature on boosting LLM performance by CoT Wei et al. (2022). Meanwhile, another key observation is the results show a certain degree of invariability to repeat numbers. Specifically, the results across different numbers of attempts don’t seem to change under either the basic or CoT strategy. Given that more repeats result in higher time and monetary cost, we chose 20 repeats for further evaluation.
Fixing the attempts to 20, Table 2 reports results by different GPT models. GPT3-turbo, GPT4-turbo, and GPT4o here point to gpt-3.5-turbo-0125, gpt-4-turbo-2024-04-09, and gpt-4o-2024-11-20. They are the most representative models of their respective series, and their release dates are roughly six months apart from one another, which may reveal invariant features of LLMs in such a prediction task.
@@
One key observation from Table 2 is that prediction performance steadily improves as LLMs iterate, though the rate of improvement decreases over time. According to Table 2, the best result under 20 repeats is achieved by GPT4o under CoT prompt strategy, which is a conclusion accuracy of 76% and a paper accuracy of 77%. Specifically, on average, GPT4o is able to predict a conclusion in 76% of the 20 repeated prediction attempts, and it also predicts 77% of the outcomes correctly for each paper. While the conclusion accuracy of GPT4o under CoT strategy is 15% higher than the CoT result of GPT3-turbo, it is only 4% higher than the CoT result of GPT4-turbo, indicating the improvement from model iteration becomes harder for this task.
@@
Another interesting finding from Table 2 is the boosting effect of CoT on performance varies on models for the experiment conclusion prediction task. Specifically, the largest improvement is around 20% on GPT4-turbo, while the least improvement is around 4% on GPT4o. However, it’s unsafe to conclude that CoT boosting is weaker on newer models since CoT improves the accuracy by 10% on the oldest model tested, GPT3-turbo.
@@
5.2 Skewness in Prediction Results
Figure 8: Conclusion Accuracy Distribution (GPT4o)
Figure 9: Paper Accuracy Distribution (GPT4o)
Two histograms (Figure 8 and 9) show the distribution of conclusion accuracy and paper accuracy obtained by GPT4o CoT prediction with 20 repeats, which is the best available result from Table 2.
As shown in Figure 8, the distribution of prediction accuracy exhibits a pronounced skewness, with a significant concentration of samples in the 90%-100% accuracy range (1165 conclusions) and another notable cluster in the 0%-10% accuracy range (278 conclusions). This indicates that the model’s performance is highly polarized, where certain conclusions are predicted with near-perfect accuracy, while others are almost wrongly predicted entirely. The similar pattern is also obvious in Figure 9 as the shape of distribution indicates that the framework performs well in a significant portion of papers.
Inspired by these unbalanced results, we closely examined the topics of each conclusion based on its experimental context. Leveraging LLMs’ ability in annotating Tan et al. (2024), we prompted Claude to label topic components of each conclusion under the context of the experiment. As a result, each conclusion got a binary vector containing only 0s and 1s like one-hot encoding vector, indicating the most related topic of it. Those topics include gender difference, socioeconomic status, ethnicity and social norms. Simiarly, Claude also decided the sentiment of each conclusion, either positive, negative, or neutral.
We find that the LLM-based prediction of experimental conclusions is sensitive to both topics and sentiments toward topics. For instance, when the conclusion is related to gender, and its context is favorable to males or detrimental to females, the conclusion accuracy is 74%. By contrast, when the context of the conclusion is favorable to females or detrimental to males, our model can make a obviously higher accuracy at over 87%. Similarly, the prediction accuracy is lower towards ethnicity-related conclusions with a negative sentiment at 80%, compared to the nearly perfect 94% of ethnicity-related conclusions with a positive sentiment. However, the LLM underperforms in the context of social norms (74%) and socioeconomic status (78%) regardless of the sentiment. In summary, we find LLMs have inclination or bias when predicting conclusions associated with specific elements because it does not fully capture mechanisms and merely make blind guesses based on their bias. This reveals scenarios on which the LLM-based field experiment simulation has limitations.
5.3 Robustness test with Manual Screening
Given that most steps in the proposed framework are automated, concerns naturally arise regarding the validity of these automated processes. To address these concerns, we conducted three manual screenings to verify the results. First, we examined whether the extracted experiment settings (Figure 3) inadvertently include genuine conclusions. Second, we checked whether the extracted conclusions align with those reported in the original papers. Finally, we assessed whether the generated variants of the conclusions (Figure 5) match our expectations.
Repeat = 20 Basic CoT
Conclusion Accuracy Paper Accuracy Conclusion Accuracy Paper Accuracy
GPT3-turbo 61% 61% 68% 67%
GPT4-turbo 66% 66% 76% 76%
GPT4o 75% 75% 78% 78%
Refined and Dequantified Sample Size: 1261 conclusions from 276 papers
Table 3: Prediction Accuracy of Refined and Dequantified samples under 20 Repeats by GPT Models
The first and third screenings revealed no issues. However, the second screening, which examined the alignment of extracted conclusions, identified 377 conclusions as either incomprehensible or nonexistent in the original texts. Consequently, after the three screenings, 1261 conclusions and 276 papers remained and were deemed valid for our purposes, reflecting approximately 77% of the initial sample size.
Additionally, 86 out of 1261 conclusions were dequantified, as existing studies suggest that predicting the magnitude of experimental outcomes remains challenging at this stage Manning et al. (2024). Given our focus on predicting the direction of experimental conclusions rather than the magnitude, conclusions specifying precise numerical treatment effects were reformulated in a dequantified manner. For instance, a conclusion such as “Job training increases income by 30%” was revised to “Job training increases income.”
Table 3 presents the aggregated findings of 1261 conclusions drawn from 276 papers. The overall performance improves marginally by approximately 2% percentage points relative to Table 2, with GPT4o retaining the highest accuracy at 78%. All observations discussed in Section 5.1 remain valid, and the skewness noted in Section 5.2 is likewise evident. Although manual screening reveals a few minor flaws in the proposed automated framework, the core conclusions—including prediction accuracy, polarization of results, model-sensitivity in CoT processes, the effect of model iteration on outcomes, and bias constraints—persist in manually refined samples, highlighting a robust reliability of the framework.
5.4 Handling Data Memorization
Figure 10: Paper Accuracy by Year
@@ 5.4 Handling Data Memorization
Data memorization is a common concern in simulating experiments with LLMs. If the results given by the LLM are from its memory of training data instead of reasoning, the proposed idea has no instructional value as pilot testing for field experiments Horton (2023). As revealed in its documents, the training data cutoff of gpt-4o-2024-11-20 is October 2023. Therefore, the report also reports accuracy by year and the standalone result for field experiments published in 2024. Notably, field experiments published after 2024 would be less likely to be included in the training data. Figure 10 shows that the performance on papers in 2024 is higher than any other, which we attribute to the small sample size of papers in 2024. Besides, for papers published from 2000 to 2023, paper accuracy of older field experiments is slightly better than more recent papers. Anyway, such a slow downward trend followed by a sudden rise in 2024 for paper accuracy is insufficient evidence of severe data memorization that denies GPT’s ability to predict field experimental conclusions.
6 Conclusions

### v2

@@ 4.3. Prediction
In the final stage, our framework takes a field experiment’s experiment settings, conclusion, and its two variants as input and generates two parallel prediction prompts: basic prediction prompt and Chain-of-Thought (CoT) prompt, which are then prompted to GPT to get predictions by asking GPT to select one conclusion from three conclusions. We also rely on CoT as it has proven to be capable of improving the general performance of LLMs on downstream tasks (Wei et al., 2022).
@@
Specifically, the basic prediction prompt is shown in Figure 10 (In Appendix B.3). It consists of a background information section, a question section, and necessary instructions. Specifically, the background information section contains the general goal of the experiment (such as exploring the impact of job training on income), treatments (such as receiving job training or not), experiment duration (such as seven weeks), outcomes (such as income), participant information (such as people seeking jobs in New England), and experiment workflow (such as when and how training was given and outcomes were recorded). All of these were extracted from a target paper, which is the same as bullet points ”A” to ”F” from Figure 8 (In Appendix B.1).
@@
Following the background information, a question section is automatically generated and entered into the templates. Specifically, the original conclusion, its reversed variant, and its unrelated variant are shuffled and substitute the placeholders ”option 1”, ”option 2”, and ”option 3”. Meanwhile, instructions tell GPT that the prediction of conclusions is under the context of the field experimental settings, asking GPT to choose one option from the three options as it deems correct.
@@
As we are also interested in how CoT would improve such prediction, Figure 11 (In Appendix B.3) shows the CoT Prediction Prompt, which follows a similar logic as the basic prediction but integrates CoT strategies to boost the performance (Wei et al., 2022). Initially, the framework prompts the experiment settings and three options for a conclusion to GPT, instructing it to think about decisive elements that help choose the correct option. Upon receiving the decisive elements from GPT, the framework prompts GPT to make a selection among three options to get a predicted conclusion. It is worth noting that the entire process is within the same GPT session for a conclusion.
@@
Although either prompt strategy generates a prediction for a given input, LLMs are stochastic models, meaning that their responses may vary to the same prompt. To handle such randomness in experiment simulations, Leng & Yuan (Leng and Yuan, 2023) set the temperature to 0 and always get fixed responses from synthetic participants of lab experiments, which is a strategy to eliminate the stochasticity of LLMs totally. By contrast, Brand et al. (Brand et al., 2023) repeated the same prompt 300 times and used the averaged number as the result of Willingness-to-Pay from customers role-played by LLMs. Here, we take the latter approach by incorporating the stochasticity of LLM outputs since this stochasticity of LLMs is similar to how the same human participant might respond differently when presented with the same instruction (Demszky et al., 2023). Specifically, we repeat the same prediction prompt several times and calculate an average accuracy as the final result. For example, if the framework is running the basic prediction, a filled-out prompt based on Figure 10 (In Appendix B.3) will be repeated a given number of times to get a stable result. The determination of a proper repeat number will be further discussed in the Results section.
@@
Parameter-wise, no fine-tuning is involved in any stage of the proposed framework, and all parameters of OpenAI API and Anthropic API are set to default. Whereas Horton (Horton, 2023) harnessed fine-tuned LLMs in lab experimental simulation to better follow instructions, Coda-Forno et al. (Coda-Forno et al., 2024) simulated several human behaviors by LLMs without fine-tuning. The use of fine-tuned LLMs complicates the reproduction attempts since other researchers don’t have access to the same model in the existing papers (Bail, 2024). Additionally, avoiding fine-tuning LLMs saves computing resources and mitigates environmental impacts (Ray, 2023), especially when the pretraining of LLMs is enough to make them capable of downstream tasks (Liu et al., 2023).
4.4. Robustness Checks
Given that most steps in the proposed framework are automated, concerns naturally arise regarding the validity of these automated processes. To address these concerns, we conduct three manual screenings to verify the results from extraction (Section 4.1) and variant generation (Section 4.2). First, we examine whether the extracted experiment settings (Figure 8, Appendix B.1) inadvertently include genuine conclusions. Second, we check whether the extracted conclusions align with those reported in the original papers. Finally, we assess whether the generated variants of the conclusions (Figure 9, Appendix B.2) match our expectations.
The first and third screenings revealed no issues. However, the second screening, which examined the alignment of extracted conclusions, identified 377 conclusions as either incomprehensible or nonexistent in the original texts. Consequently, after the three screenings, 1261 conclusions and 276 papers remained and were deemed valid for our purposes.
Additionally, 86 out of 1261 conclusions were dequantified, as existing studies suggest that predicting the magnitude of experimental outcomes remains challenging at this stage (Manning et al., 2024). Given our focus on predicting the direction of experimental conclusions rather than the magnitude, conclusions specifying precise numerical treatment effects were reformulated in a dequantified manner. For instance, a conclusion such as ”Job training increases income by 30%” was revised to ”Job training increases income.”
5. Results
@@
In this section, we test our framework on 276 field experiments that contain a total of 1261 conclusions and discuss the performance.
@@
We use Conclusion Accuracy and Paper Accuracy to evaluate the prediction performance under different settings. As shown in Figure 2, the framework generates a prompt for each conclusion. The generated prompt either follows the template in Figure 10 or 11 (In Appendix B.3), depending on which prompt strategy (basic or CoT) the framework applies. Each prompt instructs GPT to output a predicted conclusion. If the predicted conclusion matches the true conclusion, that attempt is counted as correct. As aforementioned, the framework repeats such an attempt for a set number of times for each conclusion to get a stable result. Therefore, we define Conclusion Accuracy as the percentage of correct predictions among a set number of attempts (Equation 1). Given that a field experiment may contain multiple conclusions, Paper Accuracy is the average of all Conclusion Accuracy within a paper (Equation 2).
@@
(1) Conclusion Accuracy=Num of Correct PredictionsNum of $\text{Conclusion Accuracy}=\frac{\text{Num of Correct Predictions}}{\text{Num of Predictions}}\times 100\%$
(2) Paper $\text{Paper Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\text{Conclusion Accuracy}_{i}$
5.1. Prediction Performance Overview
Table 1. Prediction Accuracy under Different Repeats
GPT-4 Turbo Basic CoT
Conclusion Acc. Paper Acc. Conclusion Acc. Paper Acc.
Repeat = 10 66% 66% 76% 76%
Repeat = 20 66% 66% 76% 76%
Repeat = 30 65% 66% 76% 76%
Sample size: 1261 conclusions from 276 papers. In the experiment we use gpt-3.5-turbo-0125, gpt-4-turbo-2024-04-09, and gpt-4o-2024-11-20.
Table 2. Prediction Accuracy under 20 Repeats by GPT Models
Model Basic CoT
Conclusion Acc. Paper Acc. Conclusion Acc. Paper Acc.
GPT-3.5 Turbo 61% 61% 68% 67%
GPT-4 Turbo 66% 66% 76% 76%
GPT-4o 75% 75% 78% 78%
Sample size: 1261 conclusions from 276 papers. In the experiment we use gpt-3.5-turbo-0125, gpt-4-turbo-2024-04-09, and gpt-4o-2024-11-20.
@@ 5.1. Prediction Performance Overview
Table 1 reports two types of accuracy for both strategies under different repeats by GPT4-turbo. The best results for the basic strategy are obtained under 10 repeats, which are 66% for both conclusion accuracy and for paper accuracy. By comparison, the CoT results are the same 76% for both strategies and different repeats, which are generally 10% points higher than basic results and aligns with prior literature on boosting LLM performance by CoT (Wei et al., 2022). Meanwhile, another key observation is that the results show a certain degree of invariability to repeat numbers. Specifically, the results across different numbers of attempts don’t seem to change under either the basic or CoT strategy. Given that more repeats result in higher time and monetary cost, we chose 20 repeats for further evaluation.
@@
Table 2 reports results by different GPT models with 20 attempts. One key observation from Table 2 is that prediction performance steadily improves as LLMs iterate, though the rate of improvement decreases over time. According to Table 2, the best result under 20 repeats is achieved by GPT4o under CoT prompt strategy, which is a conclusion accuracy of 78% and a paper accuracy of 78%. Specifically, on average, GPT4o is able to predict a conclusion in 78% of the 20 repeated prediction attempts, and it also predicts 78% of the outcomes correctly for each paper. While the conclusion accuracy of GPT4o under CoT strategy is 10 percentage points higher than the CoT result of GPT3-turbo, it is only two percentage points higher than the CoT result of GPT4-turbo, indicating the improvement from model iteration becomes harder for this task.
@@
Another interesting finding from Table 2 is the boosting effect of CoT on performance varies on models for the experiment conclusion prediction task. Specifically, the largest improvement is 10 percentage points on GPT4-turbo, while the least improvement is three percentage points on GPT4o. However, it’s unsafe to conclude that CoT boosting is weaker on newer models since CoT improves the accuracy by seven percentage points on the oldest model tested, GPT3-turbo.
@@
5.2. Data Memorization
Figure 3. Paper Accuracy by Year.
[…]
A bar chart showing paper-level prediction accuracy grouped by publication year, used to assess potential data memorization by comparing older and newer field experiments.
@@ 5.2. Data Memorization
Data memorization is a common concern in simulating experiments with LLMs. If the results given by the LLM are from its memory of training data instead of reasoning, the proposed idea has no instructional value as pilot testing for field experiments (Horton, 2023). As revealed in its documents, the training data cutoff of gpt-4o-2024-11-20 is October 2023. Therefore, it is reasonable to assume that the model would not have seen the field experiments papers appeared in 2024. In other words, field experiments published after 2024 would be less likely to be included in the training data.
To text our framework’s robustness in light of data memorization, we split the papers by year. Figure 3 shows that the paper accuracy is actually higher on papers 2024 than any other years. This is notable because if the prediction results are driven by LLMs’ memorization of its training data, the framework’s performance in unseen papers in 2024 would be poorer, indicating data memorization of less concern to our framework.
(a) Basic Conclusion Prediction Accuracy
(b) CoT Conclusion Prediction Accuracy
Figure 4. Conclusion Prediction Accuracy Distribution
[…]
Two vertically stacked bar charts showing conclusion-level prediction accuracy for basic prompting and chain-of-thought prompting.
(a) Basic Paper Prediction Accuracy
(b) CoT Paper Prediction Accuracy
Figure 5. Paper Prediction Accuracy Distribution
[…]
Two vertically stacked bar charts showing prediction-level prediction accuracy for basic prompting and chain-of-thought prompting.
5.3. Examination of Distributions of Results
To further examine the prediction performance of our proposed framework, we plot the distributions of accuracy results reported in Table 2, based on different GPT models and two prompt strategies under 20 repeats (Figure 4(b) and 5(b)).
As shown in Figure 4(a), the conclusion accuracy of all three models (gpt-3, gpt4-turbo and gpt-4o) exhibits bimodal patterns, with peaks concentrated toward both the lower and upper extremes (a U-shaped distribution). Specifically, there are significant concentrations of samples in the 90%-100% accuracy range and another notable cluster in the 0%-10% accuracy range. This indicates that the model’s performance is highly polarized, where certain conclusions are predicted with near-perfect accuracy, while others are almost wrongly predicted entirely. Interestingly, by applying CoT strategy (Figure 4(b)), the concentration in the 0%-10% accuracy range is mitigated, while the concentration in the 90%-100% accuracy range deepens, resulting from the performance boosting by CoT. Model-wise, the U-shaped distributions for the earlier model (gpt-3-turbo) are milder compared to recent models (gpt-4-turbo and gpt-4o), probably resulting from the performance limitation of earlier models.
Similar patterns also exist in Figure 5(b) to show the paper’s accuracy. Distributions for all models are skewed. Such a skewness is more pronounced in gpt-4-turbo and gpt-4o, with samples concentrating in the upper extreme. This indicates that the framework correctly predicts all conclusions from certain tested experiments. Additionally, CoT generally increases the degree of skewness(Figure 5(b)), aligning with its boosting effect on LLMs’ performance.
Inspired by the bimodal and skewed results (Section 5.3), we closely examined the topics of each conclusion based on its experimental context. Leveraging LLMs’ ability in annotating (Tan et al., 2024), we prompted Claude to label topic components of each conclusion under the context of the experiment, as shown in Figure 12. To ensure selected topics could grasp the reason behind the abnomalities found previously (Figure 4(b) and Figure 5(b)), we included topics where LLMs’ responses are biased (Hadi et al., 2023; Abid et al., 2021; Taubenfeld et al., 2024b; Lin et al., 2024; Lucy and Bamman, 2021), including gender, ethnicity, social norms, ethical dilemmas, age, socioeconomic status, and other topics. As a result, each conclusion was represented by a vector of percentages summing to 100%, where each percentage indicated the degree to which the context was associated with a particular topic. Furthermore, as the sentiment bias also affects the generated content of LLMs on top of topics bias (Oketunji et al., 2023), we used Claude as the sentiment analysis tool to label each conclusion, deciding the sentiment of each conclusion, either positive, negative, or neutral (Figure 13, Appendix B.4). In addition to regular sentiments, for gender-related conclusions, Claude also labeled each of them with a gender favorability tag, as LLMs may favor and tend to generate pro-female content (Sun et al., 2023). Specifically, if the context of a conclusion relates to gender, Claude would further judge if the context is favorable to females or detrimental to males, the opposite, or neutral, as shown in Figure 14. Finally, all variables acquired from the labeling process are summarized in Table 4 (In Appendix C).
To study the impact of topic components and sentiments on LLMs’ performance in experimental prediction, we constructed a regression model (Equation 3). Table 3 reports six regression results by three GPT models and two prompt strategies under 20 repeats, which corresponds with the conclusion accuracy results in Table 2. Notably, the sample size, which is the number of conclusions, decreases from 1261 to 955 because the content filtering policy prevents Claude from generating labels in the labeling process (Figure 121314, Appendix B.4).
{strip}
[…]
$+\beta_{4}(\text{Ethical dilemmas})+\beta_{5}(\text{Age})+\beta_{6}(\text{Socioeconomic status})$
$+\beta_{7}(\text{Other topics})+\beta_{8}(\text{Gender}\times\text{Favorability})$
$+\beta_{9}(\text{Ethnicity}\times\text{Sentiment})+\beta_{10}(\text{Social norms}\times\text{Sentiment})$
$+\beta_{11}(\text{Ethical dilemmas}\times\text{Sentiment})+\beta_{12}(\text{Age}\times\text{Sentiment})$
$+\beta_{13}(\text{Socioeconomic status}\times\text{Sentiment})+\beta_{14}(\text{Other topics}\times\text{Sentiment})$
$+\epsilon$
Figure 6. Full specification of the regression model used to examine the effect of topic–sentiment interactions on prediction accuracy.
There are several key findings from regression results (Table 3). First, the interaction effects between certain topics and sentiments significantly affect the performance, though there is no significant evidence that topic components alone may affect LLMs’ ability in predicting experimental conclusions. Specifically, sentiment interactions with complex social issues such as ethnicity, social norms, age, and other topics are significantly positive, while sentiment interactions with ethical dilemmas is significantly negative. For example, the regression of gpt4o and basic prompt (Table 3 (1)) shows that sentiment interactions with ethnicity and other topics are significantly positive, suggesting if the context of a conclusions is positive and relates to ethnicity, LLMs tend to predict it correctly, while sentiment interaction with ethical dilemmas is significantly negative, suggesting that if the context of a conclusion is negative and relates to ethical dilemmas, LLMs tend to avoid it, resulting in a wrong prediction. Surprisingly, gender, which is a topic component widely perceived to be biased in LLMs (Lucy and Bamman, 2021; Sun et al., 2023), seems not to be a biased topic in LLM-based experiment prediction. Favorability interaction with gender is only significantly positive in gpt4o prediction under CoT, suggesting that gpt4o inclines to predict pro-female conclusions and avoid pro-male conclusions under CoT strategy only in one out of six scenarios. This is also the case for the age topic, as the coefficient of its interaction is positive only in one regression. Another interesting finding is that estimated coefficients of socioeconomic status and its interaction terms are never significant according to the regression results, suggesting that LLM prediction is unbiased to conclusions related to this topic.
Second, though the directions of estimated coefficients are stable, the significant levels of the same interaction term vary across regressions due to the refinement from model iteration and prompt strategies. Specifically, the estimated coefficients of the ethnicity’s interaction term are significant across all gpt models, suggesting that the iteration LLMs didn’t fix this bias. By comparison, the coefficients of the interaction of social norms are no longer significant since gpt4-turbo, suggesting that this bias might have been fixed.
Third, applying CoT strategy might mitigate the impact of sentiment interaction with certain topics. According to the regression results in Table 3 (1) and Table 3 (2), implementing CoT turns some significant estimated coefficients of interaction terms into nonsignificant. However, this is not the case for gpt3-turbo model, as CoT brings more significant coefficients.
6. Conclusion